OpenStudy (anonymous):

What is the axis of symmetry for the function? y = x2 + 3x - 4?

OpenStudy (anonymous):

The axis of simmetry of that kind of parabola has equation:\[x={-b \over 2a}\]

OpenStudy (anonymous):

x=-b/2a. x=-3/2

OpenStudy (anonymous):

a=1, b=3,c=-4. so -b is -3. and 2a=2*1=2

OpenStudy (anonymous):

\[\sqrt{5x}\]